# Finding the particularity of the active episode of SGR J1935+2154 during which FRB 20200428 occurred: Implication from a statistics of Fermi/GBM X-ray bursts

### v1

@@ 2.3 The cumulative distribution of the fluence/flux
with the parameter values listed in Table 3. According to the above-obtained value of $S_{b}$ (or $F_{b}$), we further ignore the CCD in the low fluence/flux range and fit the high-fluence/flux distribution by a single power law (SPL), which are also shown in the right panel of Fig. 4 by the dashed lines. This result indicates that the CCD of the burst fluence/flux can basically be described by a SPL, which is consistent with the results reported in previous literature (van der Horst et al., 2012; Collazzi et al., 2015; Lin et al., 2020b).
3 Unified scaling law for magnetar bursts
@@
with $\beta_{\rm E}\sim(0.6-1.0)$, where $S_{\mathrm{E}}$ is the maximum seismic wave amplitude which determines the magnitude of the aftershocks of the earthquake as $m=\mathrm{log}_{10}(S_{\mathrm{E}})$ (Aschwanden, 2011). Earthquakes are known to be SOC phenomena. Under the long-term action of external forces, such a SOC system can always enter into a statistically stationary state characterized by spatial and temporal scale independence, without fine-tuning the system parameters. In this state, a minor change can lead to an avalanche and a collapse of the whole system. Therefore, it can be naturally understood why the aftershocks of earthquakes can have temporal clustering and power-law features of CD. Moreover, the connection between these two characteristics can in principle be understood according to the analysis by Aschwanden & McTiernan (2010), who demonstrated that the waiting time can be distributed as
@@
if the variability of the event rate shows a spike-like $\delta$-function with $\lambda$ being the mean event rate. For the burst activity of SGR J1935+2154, we have $\lambda=58.04_{-4.05}^{+4.28}$ day-1 as obtained in Section 2.
@@
Figure 5: The $x-y$ USL of XRBs. The dashed lines give the fits of the data by Eq. (12) for parameter values listed in Table 4.
As a SOC system, the earthquake phenomena had further been found to satisfy a unified scaling law (USL) between the quantities of $\Delta{t}S_{\mathrm{E}}^{-\beta_{\rm E}}L^{d_{f}}$ and $\Delta{t}^{-\alpha_{\rm E}}p(\Delta{t})_{S,L}$, where $L$ is the fault lengths, $d_{f}$ is the fractal dimension, and $p(\Delta{t})_{S,L}$ is the distribution of waiting times $\Delta{t}$ between earthquakes occurring within range $L$ whose amplitude greater than $S_{\mathrm{E}}$ (Bak et al., 2002). Here, the scaling variable $\Delta{t}S^{-\beta_{\rm E}}L^{d_{f}}$ can be interpreted as a measure of the average number of events occurring within a time interval $\Delta{t}$ and a range $L$ with an amplitude greater than $S_{\mathrm{E}}$. Then, in view of the similarity between the earthquakes and the magnetar burst activities, it is supposed that, in principle, a similar USL could also be found from the XRBs of magnetars, which reads
@@
where $x\equiv c\Delta{t}S^{\beta}$ and $y\equiv\Delta{t}^{-\alpha}P(\leq\Delta{t})$, $c$ is coefficient, and $S$ represents fluence which can be replaced with $F$ if flux is used. The scaling variable $\Delta{t}S^{\beta}$ (or $\Delta{t}F^{\beta}$) can be interpreted as a measure of the average number of bursts occurring within a time interval $\Delta{t}$ with a fluence/flux less than $S$ (or $F$). By using the parameter values of $\alpha=0.73$ derived from the CCD of waiting times and $\beta=0.69$ (or 0.95) from the CCD of fluence/flux, we plot the distribution of the XRBs of SGR J1935+2154 in the $x-y$ plane in Fig. 5, where an appropriate value $c$ is adopted to define $x$ in order to set the break value of $x$ at $1$. As shown, the distribution of the observational data indeed all behave as described in Eq. (12), which have a constant-$y$ and decaying-$y$ part with parameters listed in Table 4. This result is insensitive to the selection of the fluence/flux limitation.
@@
$>10^{-7}$0.30 $\pm$ 0.02
Fluence $S$ $>10^{-6.5}$ 1.68 -0.33 $\pm$ 0.02
(erg cm-2) $>10^{-6}$0.28 $\pm$ 0.02
$>10^{-7}$0.27 $\pm$ 0.02
Flux $F$ $>10^{-6.5}$ 19.67 -0.24 $\pm$ 0.01
(erg cm-2 s-1) $>10^{-6}$0.30 $\pm$ 0.02
In more detail, the USL indicates that the bursts locating the constant part of $f(x)$, which corresponds to the Omori law in the earthquake case, are likely to be dependent afterbursts of some main bursts. Meanwhile, the events deviating from the Omori law can be classified as independent burst events. This classification can be basically in agreement with the physical intuition that XRBs that are separated by large energy difference or long waiting times could be independent. According to this statistical classification, we plot the inferred independent and dependent bursts in Figs. 6 and 7 by using different symbols, the specific properties of which are listed in Table 6-9. In Table 5, we list the numbers of the dependent and independent bursts for the four different episodes. It can be found that the number fraction of dependent bursts in the total bursts in the FRB episode (Apr.-May 2020) is overwhelmingly higher than those in the other three episodes. Specifically, the faction of the dependent bursts (i.e., the number fraction of the data of $x<1$) can be calculated to be $\sim$33(or $\sim$36%) for the fluence (or flux) definition of $x$. The observation windows owing these dependent bursts are plotted in Fig. 7. According to this result, it can be naturally understood why the FRB episode has the smallest mean value of the waiting times, as found from Fig. 3. The large fraction of the dependent bursts increase the density of bursts in timeline and thus shorten the apparent waiting time between the successive bursts. Following this consideration, it is suspected that FRBs or slow radio bursts could exist in the last episode (Sept. 2021), because several dependent bursts appear there.
In addition, it needs to be noticed that, due to the limited detection efficiency of the instrument and the criteria used to distinguish the bursts, a bright long burst may contain an independent burst and multiple dependent bursts. The number of dependent bursts is likely underestimated. The energy of a burst may be the sum of multiple dependent ones and may be misclassified as dependent (see left panel of Fig. 6). Therefore, using the peak flux is recommended for statistical analysis of magnetar bursts.
@@ 4 Conclusion and discussion
The distribution of the waiting time and fluence/flux of the XRBs of SGR J1935+21054 are investigated statistically in this paper, which firstly show the temporal clustering of the bursts and the power-law behavior of the distributions. These results indicate that the XRBs can belong to a SOC phenomenon, which could correspond to sudden fractures and displacements of the crust of the magnetar (i.e., starquakes; Thompson & Duncan, 1995; Cheng et al., 1996). Our statistics further show that, different from the other episodes, the episode during which FRB 20200428 occurred has an obviously smaller mean value of the waiting time, which directly indicates the particularity of this episode. In more detail, by applying the USL discovered by Bak et al. (2002) for earthquakes, it is found that the regular episodes could mainly consist of independent bursts, where as much more dependent bursts can happen in the FRB episode. Nevertheless, it should be noted that we actually cannot exactly specify the individual bursts to be dependent or independent, only by according to the statistical investigation.
@@
In any case, our statistical results indicate that the starquakes occurring in regular episodes could be generally without casual link (i.e., independent). On the contrary, in the FRB episode, independent starquakes could take place more frequently and intensively and, thus, many secondary bursts could be induced by the interactions among these primary independent bursts. This further hints that FRB emission could result from such burst interactions (e.g., the collision between different seismic/Alfven waves or different explosion outflows) and be associated with a dependent burst. Such a situation could appear in a global activity of a magnetar due to a conversion of the magnetic field from a dipolar to a multi-polar configuration (e.g., the situation considered in Yang & Zhang, 2021). Here, it is interesting to mention that a giant glitch occurred approximately 3 days before FRB 20200428, as reported recently by Ge et al. (2022). The occurrence of the glitch means the magnetar crust had just obtained a sudden increase of the angular momentum, which could be transferred from the core of the magnetar. In such a case, due to the increase of the differential rotation between the crust and the core, the stress forced on the whole crust by the internal magnetic field could be consequently strengthened to exceed the elastic yield limit. Therefore, it can be imaged that crust cracking events could take place nearly globally and, simultaneously, a multi-polar field can be formed.
@@
After the FRB 20200428 event, it has further been reported that the Gravitational Wave High-energy Electromagnetic Counterpart All-sky Monitor (GECAM) detected an XRB from SGR J1935+2154 at an arrival time consistent with a radio burst detected by CHIME (Wang et al., 2022) on 2020 October 14 UTC 19:21:39.100. The Insight–Hard X-ray Modulation Telescope (HXMT) also detected an XRB associated with a radio burst detected by the Yunnan 40 m telescope, on 2022 October 21 (Li et al., 2022b). Future works on these more XRB-FRB association events would help us to test the discovery in this paper.
@@ Appendix A The X-ray bursts classification of SGR J1935+2154
Note. — The Flux in units of ×$\times 10^{7}$ erg $\cdot$ cm-2 s-1 and Fluence in units of ×$\times 10^{7}$ erg $\cdot$ cm-2. The time range of the 1st, 2nd, 3rd and 4th episodes is May 2016, June 2016, Apr.-May 2020 and Sept. 2021 respectively (see Table 2).
Note. — The Flux in units of ×$\times 10^{7}$ erg $\cdot$ cm-2 s-1 and Fluence in units of ×$\times 10^{7}$ erg $\cdot$ cm-2. The time range of the 1st, 2nd, 3rd and 4th episodes is May 2016, June 2016, Apr.-May 2020 and Sept. 2021 respectively (see Table 2).
Note. — The Flux in units of ×$\times 10^{7}$ erg $\cdot$ cm-2 s-1 and Fluence in units of ×$\times 10^{7}$ erg $\cdot$ cm-2. The time range of the 1st, 2nd, 3rd and 4th episodes is May 2016, June 2016, Apr.-May 2020 and Sept. 2021 respectively (see Table 2).
@@
Note. — The Flux in units of ×$\times 10^{7}$ erg $\cdot$ cm-2 s-1 and Fluence in units of ×$\times 10^{7}$ erg $\cdot$ cm-2. The time range of the 1st, 2nd, 3rd and 4th episodes is May 2016, June 2016, Apr.-May 2020 and Sept. 2021 respectively (see Table 2).

### v6

@@ II.3 The cumulative distribution of the fluence/flux
with the parameter values listed in Table 3. According to the above-obtained value of $S_{b}$ (or $F_{b}$), we further ignore the CCD in the low fluence/flux range and fit the high-fluence/flux distribution by a single power law (SPL), which are also shown in the right panel of Fig. 4 by the dashed lines. This result indicates that the CCD of the burst fluence/flux can basically be described by a SPL, which is consistent with the results reported in previous literature (vanderHorst2012; Collazzi2015; Lin2020a).
III Unified scaling law for magnetar bursts
@@
with $\beta_{\rm E}\sim(0.6-1.0)$, where $S_{\mathrm{E}}$ is the maximum seismic wave amplitude which determines the magnitude of the aftershocks of the earthquake as $m=\mathrm{log}_{10}(S_{\mathrm{E}})$ (Aschwanden2011book). Earthquakes are known to be SOC phenomena. Under the long-term action of external forces, such a SOC system can always enter into a statistically stationary state characterized by spatial and temporal scale independence, without fine-tuning the system parameters. In this state, a minor change can lead to an avalanche and a collapse of the whole system. Therefore, it can be naturally understood why the aftershocks of earthquakes can have temporal clustering and power-law features of CD. Moreover, the connection between these two characteristics can in principle be understood according to the analysis by AschwandenMarkus2010, who demonstrated that the waiting time can be distributed as
@@
if the variability of the event rate shows a spike-like $\delta$-function with $\lambda$ being the mean event rate. For the burst activity of SGR J1935+2154, we have $\lambda=58.04_{-4.05}^{+4.28}$ day-1 as obtained in Section II.
@@
Figure 5: The xyx-y USL of XRBs. The dashed lines give the fits of the data by Eq. (12) for parameter values listed in Table 4.
As a SOC system, the earthquake phenomena had further been found to satisfy a unified scaling law (USL) between the quantities of $\Delta{t}S_{\mathrm{E}}^{-\beta_{\rm E}}L^{d_{f}}$ and $\Delta{t}^{-\alpha_{\rm E}}p(\Delta{t})_{S,L}$, where $L$ is the fault lengths, $d_{f}$ is the fractal dimension, and $p(\Delta{t})_{S,L}$ is the distribution of waiting times $\Delta{t}$ between earthquakes occurring within range $L$ whose amplitude greater than $S_{\mathrm{E}}$ (Bak2002). Here, the scaling variable $\Delta{t}S^{-\beta_{\rm E}}L^{d_{f}}$ can be interpreted as a measure of the average number of events occurring within a time interval $\Delta{t}$ and a range $L$ with an amplitude greater than $S_{\mathrm{E}}$. Then, in view of the similarity between the earthquakes and the magnetar burst activities, it is supposed that, in principle, a similar USL could also be found from the XRBs of magnetars, which reads
@@
where $x\equiv c\Delta{t}S^{\beta}$ and $y\equiv\Delta{t}^{-\alpha}P(\leq\Delta{t})$, $c$ is coefficient, and $S$ represents fluence which can be replaced with $F$ if flux is used. The scaling variable $\Delta{t}S^{\beta}$ (or $\Delta{t}F^{\beta}$) can be interpreted as a measure of the average number of bursts occurring within a time interval $\Delta{t}$ with a fluence/flux less than $S$ (or $F$). By using the parameter values of $\alpha=0.73$ derived from the CCD of waiting times and $\beta=0.69$ (or 0.95) from the CCD of fluence/flux, we plot the distribution of the XRBs of SGR J1935+2154 in the xyx-y plane in Fig. 5, where an appropriate value $c$ is adopted to define $x$ in order to set the break value of $x$ at $1$. As shown, the distribution of the observational data indeed all behave as described in Eq. (12), which have a constant-$y$ and decaying-$y$ part with parameters listed in Table 4. This result is insensitive to the selection of the fluence/flux limitation.
@@
$10^{-7} -$0.30 $\pm$ 0.02
Fluence $S$ $10^{-6.5}$ 1.68 -0.33 $\pm$ 0.02
(erg cm-2) $10^{-6} -$0.28 $\pm$ 0.02
$10^{-7} -$0.27 $\pm$ 0.02
Flux $F$ $10^{-6.5}$ 19.67 -0.24 $\pm$ 0.01
(erg cm-2 s-1) $10^{-6} -$0.30 $\pm$ 0.02
In more detail, the USL indicates that the bursts locating the constant part of $f(x)$, which corresponds to the Omori law in the earthquake case, are likely to be dependent afterbursts of some main bursts. Meanwhile, the events deviating from the Omori law can be classified as independent burst events. This classification can be basically in agreement with the physical intuition that XRBs that are separated by large energy difference or long waiting times could be independent. According to this statistical classification, we plot the inferred independent and dependent bursts in Figs. 6 and 7 by using different symbols, the specific properties of which are listed in Tables 6-9. In Table 5, we list the numbers of the dependent and independent bursts for the four different episodes. It can be found that the number fraction of dependent bursts in the total bursts in the FRB episode (Apr.-May 2020) is overwhelmingly higher than those in the other three episodes. Specifically, the faction of the dependent bursts (i.e., the number fraction of the data of $x<1$) can be calculated to be $\sim$33% (or $\sim$36%) for the fluence (or flux) definition of $x$. The observation windows owing these dependent bursts are plotted in Fig. 7. According to this result, it can be naturally understood why the FRB episode has the smallest mean value of the waiting times, as found from Fig. 3. The large fraction of the dependent bursts increase the density of bursts in timeline and thus shorten the apparent waiting time between the successive bursts. Following this consideration, it is suspected that FRBs or slow radio bursts could exist in the last episode (Sept. 2021), because several dependent bursts appear there.
In addition, it needs to be noticed that, due to the limited detection efficiency of the instrument and the criteria used to distinguish the bursts, a bright long burst may contain an independent burst and multiple dependent bursts. The number of dependent bursts is likely underestimated. The energy of a burst may be the sum of multiple dependent ones and may be misclassified as independent (see left panel of Fig. 6). Therefore, using the peak flux is recommended for statistical analysis of magnetar bursts.
@@ IV Conclusion and discussion
The distribution of the waiting time and fluence/flux of the XRBs of SGR J1935+2154 are investigated statistically in this paper, which firstly show the temporal clustering of the bursts and the power-law behavior of the distributions. These results indicate that the XRBs can belong to a SOC phenomenon, which could correspond to sudden fractures and displacements of the crust of the magnetar (i.e., starquakes; ThompsonDuncan1995; Cheng1996). Our statistics further show that, different from the other episodes, the episode during which FRB 20200428 occurred has an obviously smaller mean value of the waiting time, which directly indicates the particularity of this episode. In more detail, by applying the USL discovered by Bak2002 for earthquakes, it is found that the regular episodes could mainly consist of independent bursts, where as much more dependent bursts can happen in the FRB episode. Nevertheless, it should be noted that we actually cannot exactly specify the individual bursts to be dependent or independent, only based on the statistical investigation.
@@
In any case, our statistical results indicate that the starquakes occurring in regular episodes could be generally without causal link (i.e., independent). On the contrary, in the FRB episode, independent starquakes could take place more frequently and intensively and, thus, many secondary bursts could be induced by the interactions among these primary independent bursts. This further hints that FRB emission could result from such burst interactions (e.g., the collision between different seismic/Alfven waves or different explosion outflows) and be associated with a dependent burst. Such a situation could appear in a global activity of a magnetar due to a conversion of the magnetic field from a dipolar to a multi-polar configuration (e.g., the situation considered in Yang2021ApJ). Here, it is interesting to mention that a giant glitch occurred approximately 3 days before FRB 20200428, as reported recently by Ge2022arxiv. The occurrence of the glitch means the magnetar crust had just obtained a sudden increase of the angular momentum, which could be transferred from the core of the magnetar. In such a case, due to the increase of the differential rotation between the crust and the core, the stress forced on the whole crust by the internal magnetic field could be consequently strengthened to exceed the elastic yield limit. Therefore, it can be imagined that crust cracking events could take place nearly globally and, simultaneously, a multi-polar field can be formed.
@@
After the FRB 20200428 event, it has further been reported that the Gravitational Wave High-energy Electromagnetic Counterpart All-sky Monitor (GECAM) detected an XRB from SGR J1935+2154 at an arrival time consistent with a radio burst detected by CHIME (Wang2022ATel) on 2020 October 14 UTC 19:21:39.100. The Insight–Hard X-ray Modulation Telescope (HXMT) also detected an XRB associated with a radio burst detected by the Yunnan 40 m telescope, on 2022 October 21 (Li2022ATel). Future works on these more XRB-FRB association events would help us to test the discovery in this paper.
@@ Appendix A The X-ray bursts classification of SGR J1935+2154
The inferred dependent/independent bursts classified by flux/fluence are listed here. The Flux in units of $\times 10^{-7}$ erg $\cdot$ cm-2 s-1 and Fluence in units of $\times 10^{-7}$ erg $\cdot$ cm-2. The time range of the 1st, 2nd, 3rd and 4th episodes is May 2016, June 2016, Apr.-May 2020 and Sept. 2021 respectively (see Table 2). Tables 6-9 present the inferred dependent/independent bursts classified by flux/fluence.